# Assessing the Influence of Toxic and Gender Discriminatory Communication on Perceptible Diversity in OSS Projects

### 12 Mar 2024 v1

@@ Abstract.
The presence of toxic and gender-identity derogatory language in open-source software (OSS) communities has recently become a focal point for researchers. Such comments not only lead to frustration and disengagement among developers but may also influence their leave from the OSS projects. Despite ample evidence suggesting that diverse teams enhance productivity, the existence of toxic or gender identity discriminatory communications poses a significant threat to the participation of individuals from marginalized groups and, as such, may act as a barrier to fostering diversity and inclusion in OSS projects. However, there is a notable lack of research dedicated to exploring the association between gender-based toxic and derogatory language with a perceptible diversity of open-source software teams. Consequently, this study aims to investigate how such content influences the gender, ethnicity, and tenure diversity of open-source software development teams. To achieve this, we extract data from active GitHub projects, assess various project characteristics, and identify instances of toxic and gender-discriminatory language within issue/pull request comments. Using these attributes, we construct a regression model to explore how they associate with the perceptible diversity of those projects.
@@ 4.1.7. Blau Index:
Gender diversity ($Blau_{Gender}$): We limit our study only to men and women since it is not possible to identify other types of gender from user names. Prior studies (Weeraddana et al., 2023) have used GenderComputer [^1] https://github.com/tue-mdse/genderComputer, genderGuesser [^2] https://pypi.org/project/gender-guesser/ and Wiki-GenderSort [^3] https://github.com/nicolasberube/Wiki-Gendersort to infer gender from people’s names. GenderComputer takes name and location as input. We will also use a combination of these tools to infer the gender of the contributors. In addition to name-based resolution, we will use GitHub profile picture-based resolution adopted in our recent study (Sultana et al., 2023).
@@ 4.2. Dataset
We intend to conduct our study on open-source projects hosted on GitHub. We employed the GitHub search tool developed by Dabic et al. (Dabic et al., 2021) for project selection. This tool allows for selecting projects based on various criteria, such as the number of contributors, forks, commits, stars, programming language, and more. Following the recommendations of Kalliamvakou et al. (Kalliamvakou et al., 2016), we filtered projects based on four criteria. The first two criteria enable the selection of software projects with active communication among contributors, ensuring sufficient data for identifying toxic and derogatory content. The remaining criteria assist in narrowing down the search space:
@@
We will also consider only non-forked projects since forked projects have a similar project history as their main ones and thus can bias our result. We also plan to divide our dataset into four groups based on the number of developers to have an in-depth idea about the occurrence and influence of toxic and derogatory language following prior studies on gender bias (Sultana et al., 2023). We estimate approximately 50,000 projects to survive this step.
@@ 4.3.2. Identifying toxic & derogatory content
We will use ToxiCR (Sarker et al., 2023) and SGID4SE (Sultana et al., 2023) for automated identfications. Following the recommendation of prior studies (Novielli et al., 2018), we plan to validate the identified toxic comments from communication excerpts. For that purpose, we will randomly select $N$ toxic comments identified by the tool and annotate them independently by at least two researchers. The value of $N$ will be computed based on Taro Yamane’s formula (Yamane, 1973) to satisfy a 95% confidence interval and 5% error margin. Similarly, we will also conduct a validation step for the identified derogatory content. We will randomly select $N$ comments from the positive class and annotate manually to assess the reliability of the tools (i.e., whether the tools’ performance degraded and their classification can be used for our study). Since ToxiCR and SGID4SE are trained for texts written in English, we will use the langdetect (lan, [n. d.]) Python library to filter out non-English texts.
@@ 4.4.7. Acceptance/rejection of hypotheses:
ratioOfDiscriminatoryComment would be the independent variable of interest

### 14 Mar 2024 v2

@@ Abstract.
The presence of toxic and gender-identity derogatory language in open-source software (OSS) communities has recently become a focal point for researchers. Such comments not only lead to frustration and disengagement among developers but may also influence their leave from the OSS projects. Despite ample evidence suggesting that diverse teams enhance productivity, the existence of toxic or gender identity discriminatory communications poses a significant threat to the participation of individuals from marginalized groups and, as such, may act as a barrier to fostering diversity and inclusion in OSS projects. However, there is a notable lack of research dedicated to exploring the association between gender-based toxic and derogatory language with a perceptible diversity of open-source software teams. Consequently, this study aims to investigate how such content influences the gender, ethnicity, and tenure diversity of open-source software development teams. To achieve this, we extract data from active GitHub projects, assess various project characteristics, and identify instances of toxic and gender-discriminatory language within issue/pull request comments. Using these attributes, we construct regression models to explore how they associate with the perceptible diversity of those projects.
@@ 4.1.7. Blau Index:
Gender diversity ($Blau_{Gender}$): We limit our study only to men and women since it is not possible to identify other types of gender from user names. Prior study (Weeraddana et al., 2023) have used GenderComputer [^1] https://github.com/tue-mdse/genderComputer, genderGuesser [^2] https://pypi.org/project/gender-guesser/ and Wiki-GenderSort [^3] https://github.com/nicolasberube/Wiki-Gendersort to infer gender from people’s names. GenderComputer takes name and location as input. We will also use a combination of these tools to infer the gender of the contributors. In addition to name-based resolution, we will use GitHub profile picture-based resolution adopted in our recent study (Sultana et al., 2023).
@@ 4.2. Dataset
We intend to conduct our study on open-source projects hosted on GitHub. We employed the GitHub search tool developed by Dabic et al. (Dabic et al., 2021) for project selection. This tool allows for selecting projects based on various criteria, such as the number of contributors, forks, commits, stars, programming language, and more. Following the recommendations of Kalliamvakou et al. (Kalliamvakou et al., 2016), we will filter projects based on four criteria. The first two criteria enable the selection of software projects with active communication among contributors, ensuring sufficient data for identifying toxic and derogatory content. The remaining criteria assist in narrowing down the search space:
@@
We will also consider only non-forked projects since forked projects have a similar project history as their main ones and thus can bias our result. We also plan to divide our dataset into four groups based on the number of developers to have an in-depth idea about the occurrence and influence of toxic and derogatory language following prior study on gender bias (Sultana et al., 2023). We estimate approximately 50,000 projects to survive this step.
@@ 4.3.2. Identifying toxic & derogatory content
We will use ToxiCR (Sarker et al., 2023) and SGID4SE (Sultana et al., 2023) for automated identfications. Following the recommendation of prior study (Novielli et al., 2018), we plan to validate the identified toxic comments from communication excerpts. For that purpose, we will randomly select $N$ toxic comments identified by the tool and annotate them independently by at least two researchers. The value of $N$ will be computed based on Taro Yamane’s formula (Yamane, 1973) to satisfy a 95% confidence interval and 5% error margin. Similarly, we will also conduct a validation step for the identified derogatory content. We will randomly select $N$ comments from the positive class and annotate manually to assess the reliability of the tools (i.e., whether the tools’ performance degraded and their classification can be used for our study). Since ToxiCR and SGID4SE are trained for texts written in English, we will use the langdetect (lan, [n. d.]) Python library to filter out non-English texts.
@@ 4.4.7. Acceptance/rejection of hypotheses:
ratioOfDiscriminatoryComment would be the independent variable of interest .